OpenStudy (anonymous):

If x varies inversely with y and x = 8 when y = 10, find y when x = 6.

Nnesha (nnesha):

do you know the equation for inverse variation ??

OpenStudy (anonymous):

yea

OpenStudy (theopenstudyowl):

y = k/x

Nnesha (nnesha):

okay so replace x and y by their values use first one x = 8 and y = 10 and find k first

OpenStudy (theopenstudyowl):

then use k to find the answer.....

OpenStudy (theopenstudyowl):

this is like an SAT question

OpenStudy (theopenstudyowl):

lol

Nnesha (nnesha):

let us know what you get @sarahefal \(\Huge\ddot{\Large\smile}\)

OpenStudy (anonymous):

ok give me one moment

Nnesha (nnesha):

\(\Huge\ddot{\Large\smile}\) okay

OpenStudy (anonymous):

80=k

Nnesha (nnesha):

yes right

Nnesha (nnesha):

now use same equation and replace k by 80 and x by 6 and solve for y

OpenStudy (anonymous):

ok hold on

Nnesha (nnesha):

\(\Huge\ddot{\Large\smile}\) okay

OpenStudy (anonymous):

10?

Nnesha (nnesha):

nope

OpenStudy (anonymous):

I did y=80/8

Nnesha (nnesha):

\[\large\rm y = \frac{ k }{ x }\] k = 80 x = 6 solve for y

OpenStudy (anonymous):

ok

Nnesha (nnesha):

nope x is not 8 this time you have to use 2nd x value

Nnesha (nnesha):